OpenStudy (jessiegonzales):

MEDAL!! 5. Solve the equation or formula for the indicated variable. S=5r^2t, for t

OpenStudy (jessiegonzales):

A. t=2/5 -r B. T=25R/S C. T=R^2 - 5S d. t=s/5r^2

OpenStudy (anonymous):

You need to get the t value by itself on one side of the equation

OpenStudy (jessiegonzales):

how do i do that

OpenStudy (anonymous):

What happens if you divide both sides by 5r^2

OpenStudy (jessiegonzales):

you get 1

OpenStudy (jessiegonzales):

Answer is D